# Optimized Fish Locomotion using Design-by-Morphing and Bayesian Optimization

### 06 Mar 2026 v2

@@ Abstract
This study presents a computational framework for optimizing undulatory swimming profiles using a combination of Design-by-Morphing and Bayesian optimization strategies. The swimming profile are expressed by morphing five baseline bio-inspired profiles using Design-by-Morphing to create an exploratory design space. The optimization objective is to find the optimal swimming profile, wavelength and undulation frequency to maximize propulsive efficiency. Arbitrary Lagrangian–Eulerian formulation is employed to simulate the unsteady flow around two-dimensional undulating swimmers. The optimized swimming profiles demonstrate a marked improvement in propulsive efficiency relative to the reference anguilliform and carangiform modes. The best-performing optimized cases achieve peak efficiencies in the range of 49%–57% over a broad range of kinematic conditions, representing an overall enhancement of 16%–35% compared to reference anguilliform and carangiform modes. The improved performance is attributed to favorable surface stress distributions and enhanced energy recovery mechanisms. A detailed force decomposition reveals that the optimal swimmer minimizes resistive drag and maximizes constructive work contributions, particularly in the anterior and posterior body regions. Spatial and temporal work decomposition indicates a strategic redistribution of input and recovered energy, enhancing performance while reducing energetic cost relative to propulsive force. These findings demonstrate that morphing-based parametric design, when guided by surrogate-assisted optimization, offers a powerful framework for discovering energetically efficient swimming gaits, with significant implications for the design of autonomous underwater propulsion systems and the broader field of bio-inspired locomotion.
@@ 1 Introduction
Optimization of undulatory kinematic profiles for self-propelling systems, such as fish-like robots, has become a pivotal area of research in bio-inspired robotics and computational fluid dynamics (CFD) (Abouhussein and Peet, 2023a; Li et al., 2025; Fernandez-Feria et al., 2022; Lin et al., 2019; Cui et al., 2018). These systems aim to mimic the natural motion of aquatic organisms, generating thrust for propulsion by effectively converting undulatory motion into mechanical energy. Such systems show significant promise in various applications, including autonomous underwater vehicles (AUVs) (Rattanasiri et al., 2015; Cui et al., 2018), where energy efficiency and maneuverability are critical performance metrics (Sun et al., 2024; Yu et al., 2017). To enhance the efficiency of these systems, optimizing the swimming profile is essential, ensuring maximum propulsion while minimizing energy consumption or maximizing the propulsive forces to propel faster. Such optimization efforts are a key source of inspiration for engineering applications, including energy-efficient AUVs, flow-energy harvesting devices (Farooq et al., 2023; Salazar et al., 2018), and bio-inspired swimming microrobots for targeted drug delivery (Tabak, 2018; Bunea and Taboryski, 2020), where achieving optimal propulsion translates directly into improved performance and extended operational capability.
@@
The fish-like swimming profiles traditionally used in CFD are categorized primarily based on the mode of body deformation, which is often quantified by the non-dimensional body wavelength $\lambda^{*}$ (the ratio of body wavelength to body length) (Sfakiotakis et al., 2002; Triantafyllou et al., 2000). These include anguilliform ($\lambda^{*}\ll 1$) motion, where the entire body undergoes large-amplitude undulations (e.g., eel); subcarangiform and carangiform ($\lambda^{*}\approx 1$) modes, where deformation is progressively confined to the posterior half or third of the body (e.g., trout and mackerel, respectively); and thunniform ($\lambda^{*}\gg 1.0$) motion, characterized by minimal body deformation with oscillations largely restricted to the caudal fin (e.g., tuna) (Khalid et al., 2021b, a). These deformation-based classifications are widely adopted in CFD to assess the hydrodynamic performance, thrust production, and propulsive efficiency of bio-inspired systems. In this study, we aim to systematically optimize the swimming profile by leveraging these existing bioinspired modes via a strongly coupled fluid–structure–optimization framework to achieve enhanced propulsive performance. Over the years, numerous studies have focused on understanding the fluid-structure interaction (FSI) in bio-inspired robotic systems (Ying et al., 2022), particularly in the context of fish-like robots (Koiri et al., 2025; Khalid et al., 2016b, 2020b, 2020a). These studies have primarily concentrated on the optimization of the kinematic and geometric properties of the robotic swimmer to improve its hydrodynamic performance.
@@
We employ 2D CFD simulations to evaluate the propulsive performance of candidate swimming profiles. Although real swimming is inherently three-dimensional, two-dimensional simulations, which are essentially slices of a biological swimmer, are widely adopted in the literature as a state-of-the-art approach due to their ability to capture the essential hydrodynamic mechanisms with significantly reduced computational cost. The swimming profiles are constructed as linear combinations of five baseline geometries, enabling exploration of a high-dimensional and continuous design space. To assess the propulsive performance of candidate configurations, full-order CFD simulations are employed, capturing the underlying FSI dynamics and evaluating metrics, such as the propulsive efficiency and power consumption.
To efficiently navigate this computationally intensive design space, we integrate BO as a surrogate-assisted global optimization strategy. BO leverages probabilistic models, such as Gaussian Processes, to predict the performance landscape and iteratively sample the most promising candidates with minimal CFD evaluations. Through this integrated DbM–BO framework, we aim to identify optimal combinations of morphing parameters and swimming kinematics that maximize propulsive efficiency, thereby advancing the design of energy-efficient, self-propelling fish-like robotic systems.
@@ 2 Methodology
Figure 1: Optimization framework combining Design-by-Morphing, CFD simulation, and Bayesian Optimization to maximize propulsive efficiency.
@@ 2.1 Design-by-Morphing
The swimming profile $A(x^{*})$, representing the lateral displacement along the fish body, is optimized using a DbM technique. This approach models the swimming profile as a linear combination of multiple baseline shapes to explore a wide range of design spaces, as illustrated in Fig. 3 for five baseline shapes. By varying the coefficients associated with these baseline profiles, we generate new swimming kinematics that aim to maximize propulsive efficiency. We considered five baseline shapes, including commonly used swimming profiles, such as anguilliform and carangiform (Khalid et al., 2021a; Maertens et al., 2016), along with the unconventional (or weird) shapes introduced to enhance diversity in the design space. These additional shapes, which differ significantly from natural swimming patterns, expand the design space and allow for a broader exploration of potential swimming profiles. Furthermore, to ensure the independence of the swimming profiles, polynomials of varying degrees are employed, except that both the anguilliform and carangiform profiles are represented using second-degree polynomials. Table 1 presents the mathematical expressions for the baseline shapes, whereas the swimming amplitudes and kinematic description of the NACA-0012 airfoil in swimming motion are depicted in Fig. 2. The swimming profile based on the design-by-morphing technique is mathematically defined as:
@@
where $N$ represents the number of baseline shapes, which in our case is five. The term $\gamma$ is the normalization factor, defined as $\gamma=\max_{i}{\left|\sum_{j=1}^{N}\omega_{j}A_{j}(x^{*}_{i})\right|}$, and $a_{n}$ denotes the maximum lateral amplitude of the swimming body, in our case, we set it as $0.1/L$, consistent to maximum tail amplitude of frequenctly used carangiform and anguilliform profiles. Furthermore, the weighting coefficients $\omega_{j}$, referred to as “weights” or “design variables” satisfy the condition $\sum_{j}\omega_{j}^{2}=1$. These weights serve as independent parameters that define the search space for optimizing the swimming profile. Enforcing the summation condition on the weights reduces the degrees of freedom by one. By utilizing a unit hypersphere centered at the origin, the five design space variables (i.e., weights) can be mapped to four hyperspherical coordinates. The transformation is expressed as:
@@
Here $\varphi_{i}$ represents the angular coordinates. Since the morphed swimming shape based on the negative weights (i.e., $\omega_{j}$) is simply the mirror image of the corresponding shape with positive weights (i.e., $\omega_{j}$), it is reasonable to consider only one-half of the hypersphere. Consequently, the angular coordinates are restricted to the range $\varphi_{i}\in\left[0,\pi\right]$.
@@ 2.2 Governing Model for Fluid Flows
Here $u_{i}$, $p$, and $\rho$ denote the $i$-th flow velocity component, pressure, and density of the fluid, respectively. The Reynolds number is defined as $\mbox{Re}={\rho LU_{\mbox{max}}}/{\mu}$, where $\mu$ is the dynamic viscosity. In this context, the spatial variables and flow velocity are non-dimensionalized using the characteristic length $L$, which is defined as the backbone of the undulating fish, and the maximum undulatory velocity $U_{\mbox{max}}$, respectively.
@@
Here $\widetilde{\mathbf{u}}=\mathbf{u}-\mathbf{u}_{\mathsf{g}}$ represents the relative flow velocity with respect to the grid node velocity, $\mathbf{u}_{\mathsf{g}}$. The computation of $\mathbf{u}_{\mathsf{g}}$ is detailed in (Farooq et al., 2022a, b). It is important to note that Eq. (8) reduces to Eq. (7) in the Eulerian description ($\widetilde{\mathbf{u}}=\mathbf{u}$), where the grid remains stationary. In the Lagrangian description ($\widetilde{\mathbf{u}}=0$), the equation tracks the variation of $\mathbf{u}$ for a fixed fluid particle.
The optimization objective in this study is defined in terms of the propulsive efficiency of the undulating foil. This measure is particularly suitable, as it reflects the balance between the useful forward propulsion and the energetic cost of the imposed motion. It is demonstrated that the definition of propulsive efficiency is controversial for propelling bodies, as the conventional thrust-based efficiency leads to a zero net thrust in the steady-state phase. Therefore, to evaluate the propulsive efficiency, Kern and Koumoutsakos Kern and Koumoutsakos (2006) proposed an alternative definition of efficiency, $\eta$, which is expressed as:
@@ 3 Numerical Verification & Validation
For the validation of the numerical solver, we first compare our results for a tethered NACA 0012 airfoil undergoing carangiform undulation at $\mbox{Re}=500$ and $\lambda^{*}=1$ over a range of undulation frequencies from 1 to 4 Hz. Our results, in terms of the time history of drag coefficient, along with those reported by Khalid et al. (2016a), is plotted in Fig. 6, showing good agreement across all considered frequencies. In addition, simulations of a tethered NACA 0012 airfoil undergoing anguilliform undulation at $\mbox{Re}=1000$ are performed over a range of undulation frequencies and non-dimensional wavelengths. The resulting time-averaged power, as shown in Fig. 6, is compared with the results of Khalid et al. (2020c), confirming the accuracy of the solver in terms of power computation.
@@
Figure 6: Validation of the present results: (a) time history of drag coefficients compared with the data of Khalid et al. (2016a), and (b) time-averaged power compared with the results of Khalid et al. (2020c).
@@ 4.1 Performance Analysis of Optimal Profiles
Name Design Epoch $\omega_{1}$ $\omega_{2}$ $\omega_{3}$ $\omega_{4}$ $\omega_{5}$ $\lambda^{*}$ $f$ $\eta$%
$opt_{1}$ 4th Best 50 -0.866 0.250 0.187 0.041 0.108 1.1 0.2 44.75
$opt_{2}$ 3rd Best 57 -0.866 0.250 0.187 -0.041 0.108 0.9 1.0 47.45
$opt_{3}$ 2nd Best 83 -0.866 0.250 0.187 -0.062 0.433 0.9 0.2 47.77
$opt_{4}$ 1st Best 105 -0.866 0.0 0.0 0.0 0.500 1.3 0.2 51.08

### 08 Apr 2026 v3

@@ Abstract
Nature has always inspired scientists and engineers to understand the underlying mechanism leading to optimal design in bio-inspired dynamics. This study presents a computational framework for optimizing undulatory swimming profiles using a combination of Design-by-Morphing and Bayesian optimization strategies. The swimming profile are expressed by morphing five baseline bio-inspired profiles using Design-by-Morphing to create an exploratory design space. The optimization objective is to find the optimal swimming profile, wavelength and undulation frequency to maximize propulsive efficiency. Arbitrary Lagrangian–Eulerian formulation is employed to simulate the unsteady flow around two-dimensional undulating swimmers. The optimized swimming profiles demonstrate a marked improvement in propulsive efficiency relative to the reference anguilliform and carangiform modes. The best-performing optimized cases achieve peak efficiencies in the range of 49%–57% over a broad range of kinematic conditions, representing an overall enhancement of 16%–35% compared to reference anguilliform and carangiform modes. The improved performance is attributed to favorable surface stress distributions and enhanced energy recovery mechanisms. A detailed force decomposition reveals that the optimal swimmer minimizes resistive drag and maximizes constructive work contributions, particularly in the anterior and posterior body regions. Spatial and temporal work decomposition indicates a strategic redistribution of input and recovered energy, enhancing performance while reducing energetic cost relative to propulsive force. These findings demonstrate that morphing-based parametric design, when guided by surrogate-assisted optimization, offers a powerful framework for discovering energetically efficient swimming gaits, with significant implications for the design of autonomous underwater propulsion systems and the broader field of bio-inspired locomotion.
@@ 1 Introduction
Optimization of undulatory kinematic profiles for self-propelling systems, such as fish-like robots, has become a pivotal area of research in bio-inspired robotics and computational fluid dynamics (CFD) (Abouhussein and Peet, 2023a; Li et al., 2025; Fernandez-Feria et al., 2022; Lin et al., 2019; Cui et al., 2018). These systems aim to mimic the natural motion of aquatic organisms, generating thrust for propulsion by effectively converting undulatory motion into mechanical energy. Such systems show significant promise in various applications, including autonomous underwater vehicles (AUVs) (Rattanasiri et al., 2015; Cui et al., 2018), where energy efficiency and maneuverability are critical performance metrics (Sun et al., 2024; Yu et al., 2017). To enhance the efficiency of these systems, optimizing the swimming profile is essential, ensuring maximum propulsion while minimizing energy consumption or maximizing the propulsive forces to propel faster. In bio-inspired systems, the propulsion efficiency is often enhanced through fin-fin interaction or fish schooling  (Khalid et al., 2018; Akhtar et al., 2007; Akhtar and Mittal, 2005; Mittal et al., 2003; Akhtar, 2003). Such optimization efforts are a key source of inspiration for engineering applications, including energy-efficient AUVs, flow-energy harvesting devices (Farooq et al., 2023; Salazar et al., 2018), and bio-inspired swimming microrobots for targeted drug delivery (Tabak, 2018; Bunea and Taboryski, 2020), where achieving optimal propulsion translates directly into improved performance and extended operational capability.
@@
The fish-like swimming profiles traditionally used in CFD are categorized primarily based on the mode of body deformation, which is often quantified by the non-dimensional body wavelength $\lambda^{*}$ (the ratio of body wavelength to body length) (Sfakiotakis et al., 2002; Triantafyllou et al., 2000). These include anguilliform ($\lambda^{*}\ll 1$) motion, where the entire body undergoes large-amplitude undulations (e.g., eel); subcarangiform and carangiform ($\lambda^{*}\approx 1$) modes, where deformation is progressively confined to the posterior half or third of the body (e.g., trout and mackerel, respectively); and thunniform ($\lambda^{*}\gg 1.0$) motion, characterized by minimal body deformation with oscillations largely restricted to the caudal fin (e.g., tuna) (Khalid et al., 2021a, b). These deformation-based classifications are widely adopted in CFD to assess the hydrodynamic performance, thrust production, and propulsive efficiency of bio-inspired systems. In this study, we aim to systematically optimize the swimming profile by leveraging these existing bioinspired modes via a strongly coupled fluid–structure–optimization framework to achieve enhanced propulsive performance. Over the years, numerous studies have focused on understanding the fluid-structure interaction (FSI) in bio-inspired robotic systems (Ying et al., 2022), particularly in the context of fish-like robots (Koiri et al., 2025; Khalid et al., 2016, 2020b, 2020a). These studies have primarily concentrated on the optimization of the kinematic and geometric properties of the robotic swimmer to improve its hydrodynamic performance.
@@
We employ 2D CFD simulations to evaluate the propulsive performance of candidate swimming profiles. Although real swimming is inherently three-dimensional, two-dimensional simulations, which are essentially slices of a biological swimmer, are widely adopted in the literature due to their ability to capture the essential hydrodynamic mechanisms with significantly reduced computational cost. Swimming profiles are constructed as linear combinations of five baseline geometries, enabling exploration of a high-dimensional continuous design space. To assess the propulsive performance of candidate configurations, full-order CFD simulations are employed, capturing the underlying FSI dynamics and evaluating metrics, such as propulsive efficiency and power consumption.
To efficiently navigate this computationally intensive design space, we integrate BO as a surrogate-assisted global optimization strategy. BO leverages probabilistic models, such as Gaussian Processes, to predict the performance landscape and iteratively sample the most promising candidates with minimal CFD evaluations. Through this integrated DbM–BO framework, our aim is to identify optimal combinations of morphing parameters and swimming kinematics that maximize propulsive efficiency, thereby advancing the design of energy-efficient, self-propelling fish-like robotic systems.
@@ 2 Methodology
Figure 1: Optimization framework combining DbM, CFD simulation, and BO to maximize propulsive efficiency.
@@ 2.1 Design-by-Morphing
The swimming profile $A(x^{*})$, representing the lateral displacement along the fish body, is optimized using a DbM technique. This approach models the swimming profile as a linear combination of multiple baseline shapes to explore a wide range of design spaces, as illustrated in Fig. 3 for five baseline shapes. By varying the coefficients associated with these baseline profiles, we generate new swimming kinematics that aim to maximize propulsive efficiency. We considered five baseline shapes, including commonly used swimming profiles, such as anguilliform and carangiform (Khalid et al., 2021b; Maertens et al., 2016), along with the unconventional (or weird) shapes introduced to enhance diversity in the design space. These additional shapes, which differ significantly from natural swimming patterns, expand the design space and allow for a broader exploration of potential swimming profiles. Furthermore, to ensure the independence of the swimming profiles, polynomials of varying degrees are employed, except that both the anguilliform and carangiform profiles are represented using second-degree polynomials. Table 1 presents the mathematical expressions for the baseline shapes, whereas the swimming amplitudes and kinematic description of the NACA-0012 airfoil in swimming motion are depicted in Fig. 2. The swimming profile based on the design-by-morphing technique is mathematically defined as:
@@
where $N$ represents the number of baseline shapes, which in our case is five. The term $\gamma$ is the normalization factor, defined as $\gamma=\max_{i}{\left|\sum_{j=1}^{N}\omega_{j}A_{j}(x^{*}_{i})\right|}$, and $a_{n}$ denotes the maximum lateral amplitude of the swimming body. In the present case, we set it as $0.1/L$, consistent to maximum tail amplitude of frequenctly used carangiform and anguilliform profiles. Furthermore, the weighting coefficients $\omega_{j}$, referred to as “weights” or “design variables” satisfy the condition $\sum_{j}\omega_{j}^{2}=1$. These weights serve as independent parameters that define the search space for optimizing the swimming profile. Enforcing the summation condition on the weights reduces the degrees of freedom by one. By utilizing a unit hypersphere centered at the origin, the five design space variables (i.e. weights) can be mapped to four hyperspherical coordinates. The transformation is expressed as:
@@
Here the $\varphi_{i}$ represents the angular coordinates. Since the morphed swimming shape based on the negative weights (i.e. $\omega_{j}$) is simply the mirror image of the corresponding shape with positive weights (i.e., $\omega_{j}$), it is reasonable to consider only one-half of the hypersphere. Consequently, angular coordinates are restricted to the range $\varphi_{i}\in\left[0,\pi\right]$.
Table 1: Mathematical expressions for the baseline shapes.
@@ 2.2 Governing Model for Fluid Flows
Here the $u_{i}$, $p$, and $\rho$ denote the $i$-th flow velocity component, pressure, and density of the fluid, respectively. The Reynolds number is defined as $\mbox{Re}={\rho LU_{\mbox{max}}}/{\mu}$, where $\mu$ is the dynamic viscosity. In this context, the spatial variables and flow velocity are non-dimensionalized using the characteristic length $L$, which is defined as the backbone of the undulating fish, and the maximum undulatory velocity $U_{\mbox{max}}$, respectively.
@@
Here, $\widetilde{\mathbf{u}}=\mathbf{u}-\mathbf{u}_{\mathsf{g}}$ represents the relative flow velocity with respect to the grid node velocity, $\mathbf{u}_{\mathsf{g}}$. The computation of $\mathbf{u}_{\mathsf{g}}$ is detailed in (Farooq et al., 2022a, b). It is important to note that Eq. (8) reduces to Eq. (7) in the Eulerian description ($\widetilde{\mathbf{u}}=\mathbf{u}$), where the grid remains stationary. In the Lagrangian description ($\widetilde{\mathbf{u}}=0$), the equation tracks the variation of $\mathbf{u}$ for a fixed fluid particle.
The optimization objective in this study is defined in terms of the propulsive efficiency of the undulating foil. This measure is particularly suitable, as it reflects the balance between the useful forward propulsion and the energetic cost of the imposed motion. It is demonstrated that the definition of propulsive efficiency is controversial for propelling bodies, as conventional thrust-based efficiency leads to zero net thrust in the steady-state phase. Therefore, to evaluate propulsive efficiency, Kern and Koumoutsakos (2006) proposed an alternative definition of efficiency, $\eta$, which is expressed as:
@@ 3 Numerical Verification & Validation
For the validation of the numerical solver, we first compare our results for a tethered NACA 0012 airfoil undergoing carangiform undulation at $\mbox{Re}=500$ and $\lambda^{*}=1$ over a range of undulation frequencies from 1 to 4 Hz. Our results, in terms of the time history of drag coefficient, along with those reported by Khalid et al. (2016), is plotted in Fig. 6, showing good agreement across all considered frequencies. In addition, simulations of a tethered NACA 0012 airfoil undergoing anguilliform undulation at $\mbox{Re}=1000$ are performed over a range of undulation frequencies and non-dimensional wavelengths. The resulting time-averaged power, as shown in Fig. 6, is compared with the results of Khalid et al. (2020c), confirming the accuracy of the solver in terms of power computation.
@@
Figure 6: Validation of the present results: (a) time history of drag coefficients compared with the data of Khalid et al. (2016), and (b) time-averaged power compared with the results of Khalid et al. (2020c).
@@ 4.1 Performance Analysis of Optimal Profiles
To further elucidate the flow–body interaction mechanisms underlying the enhanced performance of the optimal profile, instantaneous vorticity contours for the optimal, anguilliform, and carangiform swimming modes are presented at four representative time instants ($t/T=0,10,20,$ and $30$) in Fig. 16. The snapshots reveal clear differences in wake topology and vortex evolution among the three cases. In particular, the optimal profile generates a more coherent and spatially organized reverse von Kármán vortex street, characterized by stronger and more consistently shed vortical structures. This organized wake indicates more effective momentum transfer to the fluid, resulting in higher thrust production.
These flow characteristics are consistent with the trends observed in Fig. 15, where the optimal profile achieves higher propulsive speed at the cost of moderately increased work consumption. Importantly, the enhanced coherence and strength of the wake structures indicate that the additional input work is more effectively converted into useful propulsion. As a result, the relative increase in propulsive speed slightly exceeds the corresponding increase in work, leading to a more favorable performance trade-off. This fluid-dynamic evidence further substantiates that the optimal profile attains superior propulsive efficiency compared to the anguilliform profile.
@@
Figure 16: Instantaneous vorticity contours for the optimal, anguilliform, and carangiform swimming modes at four time instants ($t/T=0,10,20,$ and $30$).